# Trapping of electromagnetic waves by two dimensional plasma-metamaterial composite structures

### v1

@@ Abstract
In this work we (i) extend previous 1D studies of electromagnetic (EM) wave propagation in an over-dense plasma-metamaterial composite into two spatial dimensions and (ii) study trapping of EM waves by the composite 2D structures (barriers). Such barriers are formed when metamaterial spatially co-exists with a plasma density depletion in a form of a slab or two-dimensional density rectangular depletions (DRDs). This is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas, studied before. We find that plasma-metamaterial composite allows to trap EM waves by both slab and DRD configurations, thus forming a standing wave at the edge of an opaque region. The standing wave subsequently damps which offers applications such as heat deposition or substrate materials (micro)machining depending on EM wave intensity. The established results may find future applications such as: more efficient plasma vapour deposition, controlling EM wave propagation (EM wave trapping) in invisibility cloaks and alike. The EM wave trapping conditions are elucidated by a set of particle-in-cell (PIC) numerical simulations.
@@ I Introduction
Ref. [1] gives a good background and lays a foundation to the considered research. It is well-known that industrial applications such as vapor deposition, material surface treatment and many others use plasma as an agent or a means of delivery [2]. Generally, physical vapor deposition (PVD) is a method in-vacuum deposition that is used to produce thin films and coatings. PVD’s main feature is that material changing from a condensed phase to a vapor phase and then back to a thin film condensed phase. Common PVD processes are sputtering and evaporation. PVD is used in the manufacture of items which require thin films for mechanical, optical, chemical or electronic functions. Examples include semiconductor devices such as thin film solar panels [3], aluminized PET film for food packaging and balloons [4], and titanium nitride coated cutting tools for metalworking. Besides PVD tools for fabrication, special smaller tools (mainly for scientific purposes) have been developed [5]. Inside the Plasma Spray-Physical Vapor Deposition, or PS-PVD, ceramic powder is introduced into the plasma flame, which vaporizes it and then condenses it on the (cooler) workpiece to form the ceramic coating. Some examples of PVD include (i) Pulsed laser deposition: In which a high-power laser ablates material from the target into a vapor; (ii) Sputter deposition: In which a glow plasma discharge (usually localized around the "target" by a magnet) bombards the material sputtering some away as a vapor for subsequent deposition; (iii) Pulsed electron deposition: In which a highly energetic pulsed electron beam ablates material from the target generating a plasma stream under non-equilibrium conditions. PS-PVD, is a unique technology that enables highly tailorable functional films and coatings with various rare metal elements to be processed. This technology bridges the gap between conventional thermal spray and vapor deposition and provides a variety of coating microstructures composed of vapor, liquid, and solid deposition units.
@@
A common method of creation of plasma in these applications is gas irradiation by microwaves. Microwaves are a form of electromagnetic radiation with frequency range between 1 and 100 GHz (wavelengths between 300 and 3 mm). When a microwave enters a vessel with the plasma discharge, and the electron density $n_{e}$ is larger than the cut-off density for the incident microwave, the latter cannot propagate through plasma and is confined to its surface [6]. In microwaves, the surface wave, with a typical width of electron inertial length $c/\omega_{\rm pe}$, sustains the over-dense plasma, but the width of the generated plasma is narrow. In order to increase penetration of microwaves into the over-dense plasma, DC magnets or magnetic field coils are used. The magnetized plasma has the propagation band under the electron cyclotron frequency, and this enables to optimize this frequency by the strength of the external magnetic field. This propagation mode is the whistler mode, and a wavenumber becomes infinity when the incident wave frequency approaches the electron cyclotron frequency. The electron cyclotron resonance (ECR) heating is useful for the efficient power injection into the overdense plasma, but it is difficult to prepare a large-scale DC magnet for the plasma generation systems. Metamaterials are artificial composites made of unit patterns whose size is much smaller than the wavelength of the corresponding waves. Metamaterials give the extraordinary propagation of EM waves and it cannot occur in natural materials. Refractive index $N$ becomes negative in metamaterials because both permittivity $\varepsilon$ and permeability $\mu$ become negative [7]. The experimental verification was performed with the combination of the split-ring resonators (SRRs) as negative-$\mu$ metamaterials [8]. Theoretical and numerical simulation studies of wave propagation in negative $N$ are numerous [9, 10, 11].
@@
Refs. [11, 12] reported the generation of the second harmonic waves in negative $N$ with quadratic nonlinear response. Also the nonlinear frequency conversion process was demonstrated to be analogous to non-linear optics [13]. Also, experimental studies were performed with the combination of the arrays of metal wires as a negative-$\varepsilon$ material [14, 15, 16]. The array of metal wires has a cut-off frequency and acquires negative $\varepsilon$ under the cut-off frequency as in the over-dense plasma [17]. A composite of an overdense plasma and a negative-$\mu$ metamaterial has been investigated experimentally [18, 19, 20] and by means of numerical simulations [21, 22]. It was demonstrated that that refractive index $N$ was negative, the second harmonic wave was present in Ref. [19], and SRRs were electrically connected with plasma at a microscopic level [20], wave propagation was not clearly confirmed in the composite of an overdense plasma and a negative-$\mu$ metamaterial. Ref.[1] conducted 1D electromagnetic PIC simulations of EM wave propagation in the composite of overdense plasma and the negative-$\mu$ metamaterial. In their simulation EM waves enter the plasma-metamaterial composite, and propagate in it with a negative phase velocity and a positive group velocity. This way, Ref.[1] confirmed that the plasma-metamaterial composite has a negative refractive index. Ref. [23] gives an uptodate background to the experimental implementation of the considered research.
@@
Although, Ref.[1] conducted fully kinetic electromagnetic PIC simulations of EM wave propagation in the composite of overdense plasma and the negative-$\mu$ metamaterial, they used PIC code modified from KEMPO1 which is spatially 1 dimensional (1D). There is a clear need to extend model of Ref. [1] beyond simple 1D case. This has very strong industrial applications in plasma processing. i.e. use of negative-$\mu$ metamaterial (the plasma-metamaterial composite) will enable to increase plasma density which is crucial for plasma processing efficiency. It can be also related to modern applications such as ion thrusters and plasma based acceleration, intense X-Ray sources etc. Typical microwaves used today have frequencies in the few GHz range and the cut-off density that corresponds to the cut-off frequency $f_{\rm pe}=\omega_{\rm pe}/2\pi$ is few 1016 m-3. If this density can be increased then plasma processing will become more efficient, as the reaction rate is usually proportional to the product of number densities of the reactant species. Thus, in this work propose to extend previous 1D results [1] into two spatial dimensions and use the modified 2D (and in the future 3D) EPOCH PIC code [24]. The key novelty is that we propose to use much more general, fully kinetic regime using EPOCH code, while all previous research [21, 22] used FTDT method to simulate plasmas, hence kinetic effects of a plasma and non-linearity were not considered. These effects must be included when a high-power EM wave enters the overdense plasma under a negative-$\mu$ state.
@@
We close the introduction by stating the aims of this work that are two-fold: (i) to extend previous one dimensional studies of electromagnetic (EM) wave propagation on an over-dense plasma-metamaterial [1] composite into two spatial dimensions and (ii) to study trapping of EM waves by the composite barriers in analogy with their trapping in preformed density cavities in near-critical density plasmas [29]. We therefore propose to use modified 2D EPOCH code, as described in section 2.
@@ II The model
where where $\omega_{\rm m}$ is the magnetic resonance frequency. The modification of EPOCH code is insertion of $\mathbf{J_{m}}$. The metamaterial effect, based on $\mathbf{J_{m}}$, does not affect both electric field, $\mathbf{E}$, and plasma in this simulation. $\mathbf{J_{m}}$ is determined only by $\mathbf{B}$ and the constant value of $\omega_{\rm m}$. Such approach has proven to be very effective in modeling of bulk effects of negative-$\mu$ metamaterial (see Ref.[1] for details). The plane, linearly-polarized EM wave is excited by the external current $\mathbf{J_{s}}$ at $x=1\Delta$, the first grid point of the simulation domain, where $\Delta=r_{D}=\sqrt{\varepsilon_{0}k_{B}T_{e}/(n_{e}q_{e}^{2})}$ is the Debye length. $\mathbf{J_{s}}$ oscillates in vertical $z$-direction $\mathbf{J_{s}}=\hat{z}J_{z}=J_{0}\sin(\omega_{0}t)$. EM wave travels in $x$-direction and has components $E_{z}$ and $B_{y}$ as it travels initially in free-space. The plasma-metamaterial composite exists in the region where $25.612.8<x<$ 25.6, see Figure 2(a), that we call plasma region. Note that spatial dimensions are normalized by the EM wave inertial length $c/\omega_{0}$. When EM wave enters the plasma region it excites a different current $\mathbf{J_{p}}$ (subscript p stands for plasma), which is prescribed by the following equations:
@@
$\frac{d\gamma v_{\alpha}}{dt}=\frac{q_{\alpha}}{m_{\alpha}}\left(\mathbf{E}+\mathbf{v_{\alpha}}\times\mathbf{B}\right),$ (4)
@@
where $\mathbf{J_{p}}$ in equation (3) is different from the initial $\mathbf{J_{s}}$ because when EM wave enters plasma-metamaterial composite, electrons and ions start to oscillate and $\mathbf{J_{p}}$ is now prescribed by dielectric response of plasma and velocity of oscillating plasma particles of species $\alpha$, $\mathbf{v_{\alpha}}$. Note that when EM wave amplitude and/or plasma number density are small, plasma has the simple dispersive back-reaction to EM wave’s $\mathbf{E}$ as $\partial\mathbf{J_{p}}/\partial t\simeq\omega_{\rm pe}^{2}\mathbf{E}$ where $\omega_{\rm pe}$ is the plasma frequency. When metamaterial is present its effect is felt my resonant interaction of EM wave’s magnetic field $\mathbf{B}$. The effect of magnetic resonance can be described by appearance of another frequency $\omega_{\rm m}$ , which appears in equation (2). Note that in equation (2), the term $\partial\mathbf{J_{m}}/\partial t=\omega_{\rm m}^{2}\mathbf{B}$ is mathematically similar to $\partial\mathbf{J_{p}}/\partial t\simeq\omega_{\rm pe}^{2}\mathbf{E}$. This effectively means that presence of metamaterial is manifest by additional current $\mathbf{J_{m}}$. In this formulation, the metamaterial does not affect $\mathbf{E}$ and plasma in this simulation. $\mathbf{J_{m}}$ is determined only by $\mathbf{B}$ and the constant $\omega_{\rm m}$. Note that plasma current $\mathbf{J_{p}}$ is prescribed by nonlinear and kinetic motions of the particles and consists of the $x$- and $z$- axis components. The magnetic current density $\mathbf{J_{m}}$ is parallel to $\mathbf{B}$ and consists of the y-axis component. Therefore effective permittivity of the plasma $\varepsilon_{p}$ and the effective permeability of the metamaterial $\mu_{m}$ are expressed as $\varepsilon_{p}=1-\omega_{\rm pe}^{2}/\omega_{0}^{2}$ and $\mu_{m}=1-\omega_{\rm m}^{2}/\omega_{0}^{2}$. This implies that is EM wave’s frequency $\omega_{0}$ is smaller than both $\omega_{\rm pe}$ and $\omega_{\rm m}$ ($\omega_{0}<\omega_{\rm pe}$ and $\omega_{0}<\omega_{\rm m}$) thus both $\varepsilon_{p}$ and $\mu_{m}$ are negative. Therefore, refractive index $N=\sqrt{\varepsilon_{p}\mu_{m}}=i^{2}\sqrt{|\varepsilon_{p}||\mu_{m}|}=-\sqrt{|\varepsilon_{p}||\mu_{m}|}$ becomes negative with all of the above described consequences such a EM wave’s negative phase speed propagation in an over-dense plasma-metamaterial composite region [1].
@@
The numerical recipe how to add the additional term responsible for the presence of the metamaterials is given in appendix A. Here we focus on the details of the numerical simulation set up. We aim using periodic boundary conditions throughout this work, which means that the solution that is generated at $x=1\Delta$, the first grid cell, that travels to the left, i.e. negative $x$-direction, and re-enters the simulation domain from the far right side of the domain $x=x_{\rm max}$, should not collide with the main studied numerical solution that travels in positive $x$-direction. This is achieved by the setting $x=x_{\rm max}=8192\Delta$. It can be seen from Fig.1(a) that the left- and right- going numerical solutions never collide. The above defined notation for the Debye length can be also written as $\Delta=\lambda_{D}=v_{\rm th,e}/\omega_{\rm pe}=\sqrt{k_{B}T_{\rm e,max}/m_{\rm e}}/\omega_{\rm pe}$. Maximal temperature is set as $T_{\rm e,max}=m_{e}(0.0375c)^{2}/k_{B}$. We use the term maximal temperature because it actually varies across x-axis. This is because as in Ref.[30] we keep pressure balance by making $p_{\rm e,i}=n_{\rm e,i}(x)kT_{\rm e,i}(x)=const$, i.e. $T_{\rm e,i}(x)=T_{\rm e,i,max}[n_{0}/n_{\rm e,i}(x)]\propto 1/n_{\rm e,i}(x)$. This ensures that the total initial pressure balance $p_{\rm e,i}=const$ is fulfilled. $n_{0}$ is the background plasma number density, which is the same both for electrons and protons. The proton to electron mass ratio is set to the realistic value of 1836. The number densities $n_{\rm e,i}$ and grid sizes are varied in different numerical runs as specified in Table 1.

### v2

# Two dimensional modelling of the interaction between electromagnetic waves and plasma-metamaterial composite structures using the particle-in-cell method

@@ Abstract
In this work we (i) extend previous 1D studies of electromagnetic (EM) wave propagation in an over-dense plasma-metamaterial composite into two spatial dimensions and (ii) study blocking of EM waves by the composite 2D structures (barriers). Such barriers are formed when metamaterial spatially co-exists with a plasma density depletion in a form of a slab or two-dimensional density rectangular depletions (DRDs). This is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas, studied before. We find that plasma-metamaterial composite allows to block EM waves by both slab and DRD configurations, thus forming a standing wave at the edge of an opaque region. The standing wave subsequently damps which offers applications such as heat deposition or substrate materials (micro)machining depending on EM wave intensity. The established results may find future applications such as: more efficient plasma vapour deposition, controlling EM wave propagation (EM wave blocking) in invisibility cloaks and alike. The EM wave blocking conditions are elucidated by a set of particle-in-cell (PIC) numerical simulations.
@@ I Introduction
Ref. [1] gives a good background and lays a foundation to the considered research. It is well-known that industrial applications such as physical vapor deposition (PVD), material surface treatment and many others use plasma as an agent or a means of delivery [2]. This technology bridges the gap between conventional thermal spray and vapor deposition and provides a variety of coating microstructures composed of vapor, liquid, and solid deposition units.
@@
A common method of creation of plasma in these applications is gas irradiation by microwaves. Microwaves are a form of electromagnetic radiation with frequency range between 1 and 100 GHz (wavelengths between 300 and 3 mm). When a microwave enters a vessel with the plasma discharge, and the electron density $n_{e}$ is larger than the cut-off density for the incident microwave, the latter cannot propagate through plasma and is confined to its surface [3]. In microwaves, the surface wave, with a typical width of electron inertial length $c/\omega_{\rm pe}$, sustains the over-dense plasma, but the width of the generated plasma is narrow. Here the term over-dense plasma has the following meaning: Plasma is only transparent to EM wave with a frequency $\omega_{0}$ larger than the electron plasma frequency $\omega_{\rm pe}=\sqrt{n_{e}e^{2}/(m_{e}\varepsilon_{0})}$, where $n_{e}$ is the electron number density. In such case permittivity $\varepsilon_{p}=1-\omega_{\rm pe}^{2}/\omega_{0}^{2}$ is a positive, real number. In the opposite case when $\omega_{0}<\omega_{\rm pe}$ the permittivity $\varepsilon_{p}$ becomes negative and plasma is opaque to such EM wave. In order to increase penetration of microwaves into the over-dense plasma, DC magnets or magnetic field coils are used. The magnetized plasma has the propagation band under the electron cyclotron frequency, and this enables to optimize this frequency by the strength of the external magnetic field. This propagation mode is the whistler mode, and a wavenumber becomes infinity when the incident wave frequency approaches the electron cyclotron frequency. The electron cyclotron resonance (ECR) heating is useful for the efficient power injection into the overdense plasma, but it is difficult to prepare a large-scale DC magnet for the plasma generation systems. Metamaterials are artificial composites made of unit patterns whose size is much smaller than the wavelength of the corresponding waves. Metamaterials give the extraordinary propagation of EM waves and it cannot occur in natural materials. Refractive index $N$ becomes negative in metamaterials because both permittivity $\varepsilon$ and permeability $\mu$ become negative [4]. The experimental verification was performed with the combination of the split-ring resonators (SRRs) as negative-$\mu$ metamaterials [5]. Theoretical and numerical simulation studies of wave propagation in negative $N$ are numerous [6, 7, 8].
@@
Refs. [8, 9] reported the generation of the second harmonic waves in negative $N$ with quadratic nonlinear response. Also the nonlinear frequency conversion process was demonstrated to be analogous to non-linear optics [10]. Also, experimental studies were performed with the combination of the arrays of metal wires as a negative-$\varepsilon$ material [11, 12, 13]. The array of metal wires has a cut-off frequency and acquires negative $\varepsilon$ under the cut-off frequency as in the over-dense plasma [14]. A composite of an overdense plasma and a negative-$\mu$ metamaterial has been investigated experimentally [15, 16, 17] and by means of numerical simulations [18, 19]. It was demonstrated that that refractive index $N$ was negative, the second harmonic wave was present in Ref. [16], and SRRs were electrically connected with plasma at a microscopic level [17], wave propagation was not clearly confirmed in the composite of an overdense plasma and a negative-$\mu$ metamaterial. Ref.[1] conducted 1D electromagnetic PIC simulations of EM wave propagation in the composite of overdense plasma and the negative-$\mu$ metamaterial. In their simulation EM waves enter the plasma-metamaterial composite, and propagate in it with a negative phase velocity and a positive group velocity. This way, Ref.[1] confirmed that the plasma-metamaterial composite has a negative refractive index. Ref. [20] gives an uptodate background to the experimental implementation of the considered research. In particular Iwai et al. [20] conducted measurements of properties of transmitted microwaves and plasma parameters in a composite of double-split-ring resonators (DSRRs), which make magnetic permeability negative. In their experiments plasma was created using a 2.45 GHz microwave. They launched the microwaves with power in the range of $100-200$ W, modulated by a pulse wave at a low frequency and a low duty ratio to detect time evolutions of plasma parameters. This way, Iwai et al. [20] performed the measurements at different positions with varying distance from the supporting plate of the DSRRs, and established the enhanced wave transmission in the composite with non-uniform profiles of electron temperature. They confirmed that the non-uniformity is due to the magnetic resonant behaviour of the DSRRs at the microscopic level, while their macroscopic manifestation is making the permeability negative, enhancing wave energy transfer inside the composite.
@@
Although, Ref.[1] conducted fully kinetic electromagnetic PIC simulations of EM wave propagation in the composite of overdense plasma and the negative-$\mu$ metamaterial, they used PIC code modified from KEMPO1 which is spatially 1 dimensional (1D). There is a clear need to extend model of Ref. [1] beyond simple 1D case. This has very strong industrial applications in plasma processing. i.e. use of negative-$\mu$ metamaterial (the plasma-metamaterial composite) will enable to increase plasma density which is crucial for plasma processing efficiency. It can be also related to modern applications such as ion thrusters and plasma based acceleration, intense X-Ray sources etc. Typical microwaves used today have frequencies in the few GHz range and the cut-off density that corresponds to the cut-off frequency $f_{\rm pe}=\omega_{\rm pe}/2\pi$ is few 1016 m-3. If this density can be increased then plasma processing will become more efficient, as the reaction rate is usually proportional to the product of number densities of the reactant species. Thus, in this work propose to extend previous 1D results [1] into two spatial dimensions and use the modified 2D (and in the future 3D) EPOCH PIC code [21]. The key novelty is that we propose to use much more general, fully kinetic regime using EPOCH code, while all previous research [18, 19] used FDTD method to simulate plasmas, hence kinetic effects of a plasma and non-linearity were not considered. These effects must be included when a high-power EM wave enters the overdense plasma under a negative-$\mu$ state.
@@
We close the introduction by stating the aims of this work that are two-fold: (i) to extend previous one dimensional studies of electromagnetic (EM) wave propagation on an over-dense plasma-metamaterial [1] composite into two spatial dimensions and (ii) to study blocking of EM waves by the composite barriers in analogy with their trapping in preformed density cavities in near-critical density plasmas [26]. We therefore propose to use modified 2D EPOCH code, as described in section 2.
@@ II The model
where where $\omega_{\rm m}$ is the magnetic resonance frequency. It does not have any associated mathematical expression like the electron plasma frequency has $\omega_{\rm pe}=\sqrt{n_{e}e^{2}/(m_{e}\varepsilon_{0})}$. $\omega_{\rm m}$ is just a constant that quantifies the effect of presence metamaterial in the numerical simulation. In the experiments such as Iwai et al. [20] $\omega_{\rm m}$ is prescribed by the material properties and physical dimensions of DSRRs. The modification of EPOCH code is insertion of $\mathbf{J_{m}}$. The metamaterial effect, based on $\mathbf{J_{m}}$, does not affect both electric field, $\mathbf{E}$, and plasma in this simulation. $\mathbf{J_{m}}$ is determined only by $\mathbf{B}$ and the constant value of $\omega_{\rm m}$. According to Ref.[1] such approach has proven to be very effective in modeling of bulk effects of negative-$\mu$ metamaterial. We would to remark that to model the response of a magnetic negative (MNG) material we also use, as in [1], the Drude model, and as for the response of the epsilon negative (ENG) permittivity of plasma. However, MNG materials are usually obtained using resonant substructures such as split-ring resonators (SRR) which are generally not accurately modeled with the Drude model. In particular, perhaps some other, a more accurate model, is needed to better describe the narrow band behavior of the MNG material in the PIC simulation. This can be a subject to future investigation. The plane, linearly-polarized EM wave is excited by the external current $\mathbf{J_{s}}$ at $x=1\Delta$, the first grid point of the simulation domain, where $\Delta=r_{D}=\sqrt{\varepsilon_{0}k_{B}T_{e}/(n_{e}q_{e}^{2})}$ is the Debye length. $\mathbf{J_{s}}$ oscillates in vertical $z$-direction $\mathbf{J_{s}}=\hat{z}J_{z}=J_{0}\sin(\omega_{0}t)$. EM wave travels in $x$-direction and has components $E_{z}$ and $B_{y}$ as it travels initially in free-space. The plasma-metamaterial composite exists in the region where $25.612.8<x<$ 25.6, see Figure 2(a), that we call plasma region. The sketch of the numerical simulation is very similar to Fig.1 from Iwai et al. [1], so it is not duplicated here. Note that spatial dimensions are normalized by the EM wave inertial length $c/\omega_{0}$. When EM wave enters the plasma region it excites a different current $\mathbf{J_{p}}$ (subscript p stands for plasma), which is prescribed by the following equations:
@@
$\frac{d\left(\gamma v_{\alpha}\right)}{dt}=\frac{q_{\alpha}}{m_{\alpha}}\left(\mathbf{E}+\mathbf{v_{\alpha}}\times\mathbf{B}\right),$ (4)
@@
where $\mathbf{J_{p}}$ in equation (3) is different from the initial $\mathbf{J_{s}}$ because when EM wave enters plasma-metamaterial composite, electrons and ions start to oscillate and $\mathbf{J_{p}}$ is now prescribed by dielectric response of plasma and velocity of oscillating plasma particles of species $\alpha$, $\mathbf{v_{\alpha}}$. EPOCH is a relativistic code therefore, note that in Eq.(4), $\gamma$ is the Lorentz $\gamma$-factor. In our case it is very close to unity. Note that when EM wave amplitude and/or plasma number density are small, plasma has the simple dispersive back-reaction to EM wave’s $\mathbf{E}$ as $\partial\mathbf{J_{p}}/\partial t\simeq\omega_{\rm pe}^{2}\mathbf{E}$ where $\omega_{\rm pe}$ is the plasma frequency. When metamaterial is present its effect is felt by resonant interaction of EM wave’s magnetic field $\mathbf{B}$. The effect of magnetic resonance can be described by appearance of another frequency $\omega_{\rm m}$ , which appears in equation (2). Note that in equation (2), the term $\partial\mathbf{J_{m}}/\partial t=\omega_{\rm m}^{2}\mathbf{B}$ is mathematically similar to $\partial\mathbf{J_{p}}/\partial t\simeq\omega_{\rm pe}^{2}\mathbf{E}$. This effectively means that presence of metamaterial is manifest by additional current $\mathbf{J_{m}}$. In this formulation, the metamaterial does not affect $\mathbf{E}$ and plasma in this simulation. $\mathbf{J_{m}}$ is determined only by $\mathbf{B}$ and the constant $\omega_{\rm m}$. Note that plasma current $\mathbf{J_{p}}$ is prescribed by nonlinear and kinetic motions of the particles and consists of the $x$- and $z$- axis components. The magnetic current density $\mathbf{J_{m}}$ is parallel to $\mathbf{B}$ and consists of the y-axis component. Therefore effective permittivity of the plasma $\varepsilon_{p}$ and the effective permeability of the metamaterial $\mu_{m}$ are expressed as $\varepsilon_{p}=1-\omega_{\rm pe}^{2}/\omega_{0}^{2}$ and $\mu_{m}=1-\omega_{\rm m}^{2}/\omega_{0}^{2}$. This implies that is EM wave’s frequency $\omega_{0}$ is smaller than both $\omega_{\rm pe}$ and $\omega_{\rm m}$ ($\omega_{0}<\omega_{\rm pe}$ and $\omega_{0}<\omega_{\rm m}$) thus both $\varepsilon_{p}$ and $\mu_{m}$ are negative. Therefore, refractive index $N=\sqrt{\varepsilon_{p}\mu_{m}}=i^{2}\sqrt{|\varepsilon_{p}||\mu_{m}|}=-\sqrt{|\varepsilon_{p}||\mu_{m}|}$ becomes negative with all of the above described consequences such a EM wave’s negative phase speed propagation in an over-dense plasma-metamaterial composite region [1].
@@
The numerical recipe how to add the additional term responsible for the presence of the metamaterials is given in appendix A. Here we focus on the details of the numerical simulation set up. We use periodic boundary conditions, both in $x$- and $y$-directions, throughout this work, which numerically are the most precise of all available boundary conditions. In order to avoid the spurious reflections, and to conserve the total energy in the simulation, we do not to use open or absorbing boundary conditions in EPOCH so that to remove the wave propagating towards negative $x$ and thus reduce the computational domain. Our choice of periodic boundary conditions means that the solution that is generated at $x=1\Delta$, the first grid cell, that travels to the left, i.e. negative $x$-direction, and re-enters the simulation domain from the far right side of the domain $x=x_{\rm max}$, should not collide with the main studied numerical solution that travels in positive $x$-direction. This is achieved by the setting $x=x_{\rm max}=8192\Delta$. It can be seen from Fig.1(a) that the left- and right- going numerical solutions never collide. The above defined notation for the Debye length can be also written as $\Delta=\lambda_{D}=v_{\rm th,e}/\omega_{\rm pe}=\sqrt{k_{B}T_{\rm e,max}/m_{\rm e}}/\omega_{\rm pe}$. Maximal temperature is set as $T_{\rm e,max}=m_{e}(0.0375c)^{2}/k_{B}$. We use the term maximal temperature because it actually varies across x-axis. This is because as in Ref.[27] we keep pressure balance by making $p_{\rm e,i}=n_{\rm e,i}(x)kT_{\rm e,i}(x)=const$, i.e. $T_{\rm e,i}(x)=T_{\rm e,i,max}[n_{0}/n_{\rm e,i}(x)]\propto 1/n_{\rm e,i}(x)$. This ensures that the total initial pressure balance $p_{\rm e,i}=const$ is fulfilled. $n_{0}$ is the background plasma number density, which is the same both for electrons and protons. The proton to electron mass ratio is set to the realistic value of 1836. The number densities $n_{\rm e,i}$ and grid sizes are varied in different numerical runs as specified in Table 1.
@@
III The results
In is section we present the results of our numerical simulations detailed in table 1. Below we also provide the justification for the number density expressions used in this work.